OpenStudy (anonymous):

Barbara made a punch by mixing quarts of ginger ale with quarts lemonade. How many pints of punch did Barbara make? (1 quart = 2 pints)

OpenStudy (anonymous):

1 1/2 quarts and 2 3/4 quarts

OpenStudy (aveline):

What is the total amount of quarts that were made? After you find this number, multiply by two to find out how many pints were made.

OpenStudy (anonymous):

how do you figure this out the first number is 1 1/2 second number is 2 3/4 quartz?

OpenStudy (anonymous):

@Aveline

OpenStudy (aveline):

The easiest way to do this would be to convert the fractions to decimals. 1 1/2 = 1.5 2 3/4 = 2.75 2.75+1.5=5.25 quarts There are 2 pints for every quart Therefore, 5.25*2=10.5 pints

OpenStudy (anonymous):

the choices are 2 1/8 6 1/4 6 1/2 8 1/2

OpenStudy (anonymous):

how do you figure that out

OpenStudy (anonymous):

10.5 converted into a fraction?

OpenStudy (aveline):

I made a math error o3o 2.75+1.5 is not 5.25, it's 4.25 You multiply 4.25*2=8.5, which is 8 1/2 pints Sorry about that